OpenStudy (anonymous):

Which lines have negative slopes? Choose all answers that are correct. x = –3 5x + 2y = 10 –3x – y = 6 –2x + 4y = 16

OpenStudy (anonymous):

Put all these equation in the form: \(y = mx + h\), where \(m\) is the slope and \(h\) is the y-intercept

OpenStudy (anonymous):

after that just choose negative m

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

i'll make one of than, then u do for the others, ok? 5x+2y=10 let's put that in teh form y = mx + h: 2y = -5x + 10 y = (-5/2)x + 5 Now, we have: m = -5/2, and h = 5 The slope (m) is negative. Got it?

OpenStudy (anonymous):

oh okay thxs i got it

OpenStudy (anonymous):

No problem ;)